OpenStudy (anonymous):

Find the mean of the collection of numbers: 3, 1, 9, 5, 9, 6, 9

OpenStudy (shadeyghost):

Mean is average correct?

jimthompson5910 (jim_thompson5910):

Add up the numbers and divide that sum by 7.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the sum is 42

OpenStudy (anonymous):

and 42 divided by 7 is 6

OpenStudy (anonymous):

so omg ur answer will be 6

OpenStudy (anonymous):

ok

OpenStudy (ahsome):

Mean is the average of the set of numbers Mean can be found by using the equation: \[Mean=\frac{\sum x}{n}\] Where: \(\sum x\) is the number you get when you add up all the numbers \(n=\) is the number of itmes there In this case: \[\sum x= 3+ 1+ 9+5+ 9+ 6+ 9\]\[\sum x=42\]\[n=7\] Sub these values into the equation: \[Mean=\frac{\sum x}{n}\]\[Mean=\frac{42}{7}\]\[Mean=6\]

OpenStudy (ahsome):

That should help explain how @ade1010 got his answe, @omg1help

OpenStudy (anonymous):

thank you @Ahsome

OpenStudy (ahsome):

No problem @omg1help :)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thx for medal omg